New Products pays no dividend at the present time. The firm will pay the first dividend of $0.25 per share at the end of Year 3. After that, the dividend will growth at a rate of 15% per year for two years. The company’s stock can be sold for $38 per share at the end of Year 5. How much should you pay per share to purchase this stock today at a required return of 13.8 percent? (Note: keep four decimal places when calculating dividends.)
Answer #1

D3=0.25

D4=(0.25*1.15)=0.2875

D5=(0.2875*1.15)=0.3306

Value after year 5=$38

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=0.25/1.138^3+0.2875/1.138^4+0.3306/1.138^5+38/1.138^5

=$20.42(Approx).